OpenStudy (anonymous):

Suppose a 15-minute call costs $5.26, and a 30-minute call costs $10.24. Find a formula relating time (in minutes) and cost (in dollars). How do you do this???

OpenStudy (dtan5457):

You need to find the slope, and create a line (function) by getting it into the slope point formula. your two points are 15,5.26 30,10.24 y2-y1/x2-x1 slope=0.332x point slope formula is y-y1=m(x-x1) y-10.24=0.332(x-30) y=0.332x+0.28<<<<your function